OpenStudy (anonymous):

Look at the quadratic equation below. -x2 + 3 = (2x - 1)2 What is this quadratic equation written in standard form?

OpenStudy (anonymous):

it is not in standar form

OpenStudy (anonymous):

\[-x^2 + 3 = (2x - 1)^2 \] \[-x^2+3=4x^2-4x+1\] is a start

OpenStudy (anonymous):

\[ax^2+bx+c=0\]is what you are aiming for

OpenStudy (anonymous):

So the answer will be 5x^2-4x+4=0

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

\[5x^2-4x-2=0\] you need to subtract \(3\) from both sides

OpenStudy (anonymous):

3x^2-4x+4=0?

OpenStudy (anonymous):

I'm confused now

OpenStudy (anonymous):

what do you mean by sign or not?